The total no. Of no. Less than 1000 and divisible by 5 formed with 0,1,2.......9 such that each digit does not occur more than once in each no.is
Answer by greenestamps(13198)


If the number is divisible by 5, then the last digit is either 0 or 5. So you have two cases to analyze.

If the last digit is 0, then there are 9 choices for one of the other digits and then 8 choices for the third digit. By the fundamental counting principle, the number of numbers less than 1000, divisible by 5, ending in 0, and containing 3 different digits is 9*8 = 72.

If the last digit is 5, there are still 9 choices for one of the other digits. That is because the problem said "numbers less than 1000", not "3-digit numbers" -- so the first digit can be 0. And then again there are 8 choices for the last digit, making again 9*8=72 numbers less than 1000, divisible by 5, ending in 5, and containing 3 different digits.

So there are a total of 72+72=144 numbers that are less than 1000, are divisible by 5, and contain 3 different digits.